OpenStudy (anonymous):

Write the percent as a decimal and fraction. 505%

OpenStudy (anonymous):

this is a hard one.....or is it?

OpenStudy (anonymous):

5.05

OpenStudy (anonymous):

5.05 is the decimal what is the fraction?

OpenStudy (anonymous):

if you give me the fraction you get a quick metal!

OpenStudy (anonymous):

101/20

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

5.05 decimal converting to fraction 5.05/1=5.05/1*100/100=505/100 then simplify